# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v2

@@ IV Implementation of Privacy-Preserving DMPC
In this section, the overall DMPC implementation is described based on the differentially-private distributed dual-gradient algorithm.
@@ IV-A Algorithm Implementation
Algorithm 3 will terminate after $\bar{k}$ iterations. Note that Algorithm 3 converges almost surely in a probability sense, and thus the global constraints (3) may not be satisfied within a given number of iterations. Based on (8), one can verify that the global constraints are satisfied if the following condition holds:
@@
To verify whether the global constraints are satisfied after the termination of Algorithm 3, we employ a privacy-preserving static average consensus method developed in [34].
@@
where $\iota$, $a_{i,\alpha\beta}^{\ell}$, $a_{i,j}^{\ell}\in\mathbb{R}_{+}$. As proven in [34], by appropriately selecting the parameters $\iota$, $a_{i,\alpha\beta}^{\ell}$, and $a_{i,j}^{\ell}$, $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ converge to the average consensus value $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ (i.e., $\frac{1}{M}\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$). Therefore, each subsystem can utilize the converged value of $z_{i,\alpha}^{\ell}$ to check whether condition (38) is satisfied. It is worth noting that conventional static average consensus approaches [35, 36, 37] can also be employed to calculate the value of $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ in a distributed manner. However, these approaches necessitate subsystems to directly share $z_{i}^{0}$ with their neighbors, potentially leading to privacy breaches as $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$ contains sensitive information about $\tilde{{\bm{u}}}_{i}^{\bar{k}}$. The average consensus method developed in [34] employs a state decomposition scheme to mask the real values of $z_{i}^{0}$. As shown in (39), the substate $z_{i,\alpha}^{\ell}$ governs the role of internode interactions and is the only value from subsystem $i$ that can be seen by its neighbors. On the other hand, the other substate $z_{i,\beta}^{\ell}$ participates in the distributed interactions by solely interacting with $z_{i,\alpha}^{\ell}$. Hence, the existence of $z_{i,\beta}^{\ell}$ is invisible to neighboring nodes of subsystem $i$, although it directly affects the evolution of $z_{i,\alpha}^{\ell}$. Through this state decomposition design, strong privacy preservation can be guaranteed. For further details, please refer to [34].
After executing the static average consensus method, an update mechanism is designed for the control input sequence $\tilde{{\bm{u}}}_{i}(t)$. Based on the consensus results, if condition (38) is met, then the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ at the current time instant is applied to $\tilde{{\bm{u}}}_{i}(t)$; otherwise, the control input sequence from the last time instant, i.e., $\tilde{{\bm{u}}}_{i}(t-1)$, is used to update $\tilde{{\bm{u}}}_{i}(t)$. The overall DMPC strategy is presented in Algorithm 4. By utilizing the static average consensus method and the update mechanism designed for $\tilde{{\bm{u}}}_{i}(t)$, we can ensure that if the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ at time instant $t=0$ is feasible, then Algorithm 4 can generate feasible solutions for the remaining time.
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 3 with $x_{i}(t)$;
@@ 1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
$\tilde{{\bm{u}}}_{i}(t)=$$\{\tilde{u}_{i}(1|t-1),\tilde{u}_{i}(2|t-1),\cdots,$
$\quad\tilde{u}_{i}(N-1|t-1),K_{i}\tilde{x}_{i}(N|t-1)\};$
@@ IV-B Feasibility and Stability
At any time instant $t$, the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 is constrained in the bounded set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$, and thus $J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ is bounded and the following relation holds:
$\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{\bar{k}})-\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})\leq\eta,$ (40)
where $\eta\in\mathbb{R}_{+}$ is a bounded constant. The following theorem summarizes the theoretical results of the developed DMPC strategy.
@@
Assume that $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 satisfies the global constraints at time instant $t=0$. Then, the following results hold:
  1. 1.
    If Algorithm 4 has a feasible solution at time instant $t$, then it has a feasible solution at $t+1$.
  2. 2.
    If $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, then the state trajectory of each subsystem converges to the terminal set $\mathcal{X}^{f}_{i}$ in finite time.
@@
As shown in Algorithm 4, the input sequence at time instant $t$ is denoted by $\tilde{{\bm{u}}}_{i}(t)=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$. Let $\tilde{{\bm{x}}}_{i}(t)=\{\tilde{x}_{i}(0|t),\tilde{x}_{i}(1|t),\cdots,\tilde{x}_{i}(N|t)\}$ be the corresponding predicted state sequence. Since $\tilde{{\bm{u}}}_{i}(t)$ is a feasible solution, it can be obtained from (7), (8), and (38) that $\tilde{{\bm{u}}}_{i}(t)\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ and
@@
Based on (9), (41), and (42), it can be concluded that $\hat{{\bm{u}}}_{i}(t+1)\in\tilde{\mathcal{U}_{i}}(x_{i}(t+1))$ and
$\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(\ell|t+1)+\varPsi_{u_{i}}\hat{u}_{i}(\ell|t+1)$ (43)
$=$ $\sum_{i=1}^{M}\varPsi_{x_{i}}\tilde{x}_{i}(\ell+1|t)+\varPsi_{u_{i}}\tilde{u}_{i}(\ell+1|t)$
$\leq$ $(1-\varepsilon M(\ell+1)){\bf 1}_{p},\quad\ell\in\mathbb{Z}_{0}^{N-2},$
$\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(N-1|t+1)+\varPsi_{u_{i}}\hat{u}_{i}(N-1|t+1)$
$=$ $\sum_{i=1}^{M}(\varPsi_{x_{i}}+\varPsi_{u_{i}}K_{i})\tilde{x}_{i}(N|t)\leq(1-\varepsilon MN){\bf 1}_{p}.$
Therefore, $\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at time instant $t+1$, which completes the proof for the first statement of Theorem 4. From the above analysis, it is evident that at $t=0$, if $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 is feasible, then the update mechanism designed for $\tilde{{\bm{u}}}_{i}(t)$ in Algorithm 4 ensures the solution feasibility for the remaining duration.
To prove the second statement, we first define a Lyapunov function $V(\{x_{i}(t)\}):=\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})$. According to the algebraic Riccati equation (6) and (42), we have
$J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))$ (44)
$=$ $-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}.$
$\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at $t+1$ but may not be optimal. Thus, we have
$V(\{x_{i}(t+1)\})$ (45)
$\leq$ $\sum_{i=1}^{M}J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))$
$=$ $\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}\right)$
$\leq$ $V(\{x_{i}(t)\})+\eta-\sum_{i=1}^{M}\|x_{i}(t)\|^{2}_{Q_{i}},$
where the equality condition is due to (44) and the last inequality follows from (40). (45) indicates that $x_{i}(t)$ converges to the bounded set $\{\{x_{i}\}:\sum_{i=1}^{M}\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}$ in finite time. Considering this fact and the assumption that $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, it can be concluded that $x_{i}(t)$ enters the terminal set $x_{i}(t)$ in finite time. ∎
@@ V Numerical Simulations
In this section, simulation is conducted to demonstrate the performance of the developed method. A group of four linear time-invariant subsystems are considered. The network structure of these four subsystems is shown in Figure 1. The system matrices $A_{i}$ and $B_{i}$ are chosen as
$A_{i}$ $=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},B_{i}=\begin{bmatrix}1\\ 1\end{bmatrix},i=1,3,$
$A_{i}$ $=\begin{bmatrix}2&1\\ 0&1\end{bmatrix},B_{i}=\begin{bmatrix}1\\ 1\end{bmatrix},i=2,4.$
For all subsystems, the local state and input constraint sets are selected as $\mathcal{X}_{i}=\{x_{i}:-1\leq x_{i}\leq 1\}$ and $\mathcal{U}_{i}=\{u_{i}:-0.3\leq u_{i}\leq 0.3\}$, respectively. The global constraint is $-0.65\leq\sum_{i=1}^{4}u_{i}\leq 0.65$. The weight matrices $Q_{i}$ and $R_{i}$ are set as $Q_{i}=I$ and $R_{i}=0.1$, respectively. The length of the prediction horizon is chosen as $N=5$. In Algorithm 3, we inject Laplace noise with parameter $\nu^{k}=0.1+0.001k^{0.1}$. The weakening factor sequence and step-size sequence is set as $\chi^{k}=\frac{2}{1+0.01k^{0.9}}$ and $\gamma^{k}=\frac{5}{1+0.1k}$, respectively. In the simulation, Algorithm 4 is executed 20 times, and the mean and the variance of the state and input trajectories are computed. For comparison, we also run Algorithm 2 under the same noise level.
Figure 1: Interaction network.
The simulation results are illustrated in Figures 2 and 3. Figure 2 depicts the evolution of the system state. It can be seen that the variance of the system state trajectories under Algorithm 2 is much larger than those under Algorithm 4. This discrepancy arises from the direct integration of persistent noise into Algorithm 1, which sacrifices optimization accuracy and subsequently degrades the control performance of Algorithm 2. In our developed approach, the weakening factor $\chi^{k}$ is tailored to alleviate the influence of DP noise, ensuring accurate convergence of the system state. In addition, Figure 3 presents the evolution of the global constraint. It can be found that there exist constraint violations in Algorithm 2. However, owing to the implementation scheme developed in Section IV, our approach can guarantee the satisfaction of the global constraint.
Figure 2: System state evolution.
Figure 3: Global constraint evolution.

### v3

@@ IV Implementation of Privacy-Preserving DMPC
In this section, the overall implementation strategy of DMPC is described.
@@ IV-A Algorithm Implementation
Algorithm 3 will terminate after $\bar{k}$ iterations. Note that Algorithm 3 converges almost surely in a probabilistic sense, and thus the global constraints (2) may not necessarily be satisfied within a finite number of iterations. Based on (7), one can verify that the global constraints are satisfied if the following condition holds:
@@
To verify whether the global constraints are satisfied after the termination of Algorithm 3, we employ a privacy-preserving static average consensus method developed in [WangTAC2019].
@@
where $\iota$, $a_{i,\alpha\beta}^{\ell}$, $a_{i,j}^{\ell}\in\mathbb{R}_{+}$. As proven in [WangTAC2019], by appropriately selecting the parameters $\iota$, $a_{i,\alpha\beta}^{\ell}$, and $a_{i,j}^{\ell}$, $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ converge to the average consensus value $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ (i.e., $\frac{1}{M}\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$). Therefore, each subsystem can utilize the converged value of $z_{i,\alpha}^{\ell}$ to check whether condition (33) is satisfied. While conventional static average consensus methods [Olfati-Saber2007, SundaramACC2007, HendrickxTAC2015] can compute $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ in a distributed manner, they require subsystems to directly share $z_{i}^{0}$ with neighbors. Since $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$ contains sensitive information about $\tilde{{\bm{u}}}_{i}^{\bar{k}}$, these methods may potentially lead to privacy breaches. The privacy-preserving average consensus method in [WangTAC2019] addresses this issue using a state decomposition scheme to mask the true values of $z_{i}^{0}$. Specifically, as shown in (34), the substate $z_{i,\alpha}^{\ell}$ governs internode interactions and is the only value visible to a subsystem’s neighbors. Meanwhile, the other substate $z_{i,\beta}^{\ell}$ interacts solely with $z_{i,\alpha}^{\ell}$, remaining hidden from neighboring subsystems but still influencing $z_{i,\alpha}^{\ell}$’s evolution. This design ensures strong privacy protection. For more details, please refer to [WangTAC2019].
The overall DMPC strategy is presented in Algorithm 4. After executing the static average consensus method, an update mechanism is designed for the control input sequence $\tilde{{\bm{u}}}_{i}(t)$. Based on the consensus results, if condition (33) is met, the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ obtained at the current time instant is adopted as $\tilde{{\bm{u}}}_{i}(t)$. Otherwise, we implement a one-step time-shift on the previous control input sequence $\tilde{{\bm{u}}}_{i}(t-1)$ and append a terminal control action to update $\tilde{{\bm{u}}}_{i}(t)$, as shown in (35). The sequence constructed via (35) is guaranteed to be feasible, which will be demonstrated in the first statement of Theorem 4. Therefore, by combining the static average consensus method with the update mechanism for $\tilde{{\bm{u}}}_{i}(t)$, Algorithm 4 ensures that if the initial solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ at $t=0$ is feasible, then feasible solutions will be maintained at all subsequent time steps—even in cases when Algorithm 3 fails to generate feasible solutions at some time instants or over multiple consecutive steps.
1At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 3;
@@
$\tilde{{\bm{u}}}_{i}(t)=$ $\{\tilde{u}_{i}(1|t-1),\tilde{u}_{i}(2|t-1),\cdots,$ (35)
$\qquad\tilde{u}_{i}(N-1|t-1),K_{i}\tilde{x}_{i}(N|t-1)\};$
@@ IV-B Feasibility and Stability
Assume that $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 satisfies the global constraints at time instant $t=0$. Then, the following results hold: $\left.1\right)$ If Algorithm 4 has a feasible solution at time instant $t$, then it has a feasible solution at $t+1$; $\left.2\right)$ $\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))-\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})\leq\eta$, where $\eta\in\mathbb{R}_{+}$ is a bounded constant; $\left.3\right)$ If $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, then the state trajectory of each subsystem converges to the terminal set $\mathcal{X}^{f}_{i}$ in finite time.
@@
As shown in Algorithm 4, the input sequence at time instant $t$ is denoted by $\tilde{{\bm{u}}}_{i}(t)=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$. Let $\tilde{{\bm{x}}}_{i}(t)=\{\tilde{x}_{i}(0|t),\tilde{x}_{i}(1|t),\cdots,\tilde{x}_{i}(N|t)\}$ be the corresponding predicted state sequence. Since $\tilde{{\bm{u}}}_{i}(t)$ is a feasible solution, it can be obtained from (6), (7), and (33) that $\tilde{{\bm{u}}}_{i}(t)\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ (i.e., $\tilde{x}_{i}(\ell|t)\!\in\!\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\!\in\!\mathcal{U}_{i},\tilde{x}_{i}(N|t)\!\in\!\mathcal{X}_{i}^{f},\ell\!\in\!\mathbb{Z}_{0}^{N-1}$) and
@@
Based on (8), (36), and (37), it can be concluded that $\hat{{\bm{u}}}_{i}(t+1)\in\tilde{\mathcal{U}_{i}}(x_{i}(t+1))$, $\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(\ell|t\!+\!1)\!+\!\varPsi_{u_{i}}\hat{u}_{i}(\ell|t\!+\!1)\!=\!\sum_{i=1}^{M}\varPsi_{x_{i}}\tilde{x}_{i}(\ell\!+\!1|t)\!+\!\varPsi_{u_{i}}\tilde{u}_{i}(\ell\!+\!1|t)\leq(1-\varepsilon M(\ell+1)){\bf 1}_{p},\ell\in\mathbb{Z}_{0}^{N-2}$, and $\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(N\!-\!1|t\!+\!1)\!+\!\varPsi_{u_{i}}\hat{u}_{i}(N\!-\!1|t\!+\!1)\!=\!\sum_{i=1}^{M}(\varPsi_{x_{i}}\!+\!\varPsi_{u_{i}}K_{i})\tilde{x}_{i}(N|t)\leq(1\!-\!\varepsilon MN){\bf 1}_{p}$. Therefore, $\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at time instant $t+1$, which completes the proof for the first statement of Theorem 4. From the above analysis, it follows that the control input sequence constructed in (35) is feasible. Thus, if $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ computed by Algorithm 3 is feasible at $t=0$, then the update mechanism for $\tilde{{\bm{u}}}_{i}(t)$ in Algorithm 4 ensures the solution feasibility for the remaining duration.
Due to the recursive feasibility, $x_{i}(t)$ remains within the bounded set $\mathcal{X}_{i}$, and the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated by Algorithm 3 is confined to the bounded set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$. Thus, $J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))$ is bounded, and there exists a positive bounded constant $\eta$ such that $\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))\!-\!\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})\leq\eta$.
To prove the third statement, we first define a Lyapunov function $V(\{x_{i}(t)\}):=\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})$. According to the algebraic Riccati equation (5) and (37), we have
$J_{i}(x_{i}(t\!\!+\!\!1),\hat{{\bm{u}}}_{i}(t\!\!+\!\!1))\!-\!J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))\!=\!\!-\|x_{i}(t)\|^{2}_{Q_{i}}\!\!-\!\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}.$ (38)
$\hat{{\bm{u}}}_{i}(t\!+\!1)$ is a feasible solution at $t+1$ but may not be optimal. Thus, we have $V(\{x_{i}(t\!+\!1)\})\leq\sum_{i=1}^{M}J_{i}(x_{i}(t\!+\!1),\hat{{\bm{u}}}_{i}(t\!+\!1))=\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))\!-\!\|x_{i}(t)\|^{2}_{Q_{i}}\!-\!\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}\right)\leq\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))\!-\!\|x_{i}(t)\|^{2}_{Q_{i}}\right)$, where the equality condition is due to (38). By Statement $\left.2\right)$, we have $V(\{x_{i}(t+1)\})\leq V(\{x_{i}(t)\})+\eta-\sum_{i=1}^{M}\|x_{i}(t)\|^{2}_{Q_{i}}$, which indicates that $x_{i}(t)$ converges to the bounded set $\{\{x_{i}\}:\sum_{i=1}^{M}\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}$ in finite time. Considering the assumption that $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\!\leq\!\eta\}\!\subset\!\mathcal{X}^{f}_{i}$, it can be concluded that $x_{i}(t)$ enters the terminal set $\mathcal{X}^{f}_{i}$ in finite time. ∎
@@ V Numerical Simulations
In this section, simulation is conducted to demonstrate the performance of the developed method. A group of four linear time-invariant subsystems are considered. The interaction weight matrix $L$ is set as $L_{12}=L_{21}=L_{14}=L_{41}=\frac{1}{4}$, $L_{23}=L_{32}=\frac{3}{8}$, $L_{34}=L_{43}=\frac{5}{16}$, $L_{13}=L_{31}=L_{24}=L_{42}=0$, $L_{11}=-\frac{1}{2}$, $L_{22}=-\frac{5}{8}$, $L_{33}=-\frac{11}{16}$, and $L_{44}=-\frac{9}{16}$. The system matrices $A_{i}$ and $B_{i}$ are chosen as
$A_{i}\!=\!\begin{bmatrix}1&1\\ 0&1\end{bmatrix},B_{i}\!=\!\begin{bmatrix}1\\ 1\end{bmatrix},i\!=\!1,3;A_{i}\!=\!\begin{bmatrix}2&1\\ 0&1\end{bmatrix},B_{i}\!=\!\begin{bmatrix}1\\ 1\end{bmatrix},i\!=\!2,4.$
For all subsystems, the local state and input constraint sets are selected as $\mathcal{X}_{i}\!=\!\{x_{i}:-1\!\leq\!x_{i}\!\leq\!1\}$ and $\mathcal{U}_{i}\!=\!\{u_{i}:-0.3\!\leq\!u_{i}\!\leq\!0.3\}$, respectively. The global constraint is $0.65\!\leq\!\sum_{i=1}^{4}u_{i}\!\leq\!0.65$. The weight matrices $Q_{i}$ and $R_{i}$ are set as $Q_{i}\!=\!I$ and $R_{i}\!=\!0.1$, respectively. The length of the prediction horizon is chosen as $N=5$. In Algorithm 3, we inject Laplace noise with parameter $\nu^{k}\!=\!0.1\!+\!0.001k^{0.1}$. The weakening factor sequence and step-size sequence is set as $\chi^{k}=\frac{2}{1+0.01k^{0.9}}$ and $\gamma^{k}=\frac{5}{1+0.1k}$, respectively. In the simulation, Algorithm 4 is executed 20 times, and the mean and the variance of the state and input trajectories are computed. For comparison, we also run Algorithm 2, which uses Algorithm 1 for distributed computation, and apply the same level of noise to the shared variables in Algorithm 1.
The simulation results are illustrated in Fig. 2(a). Fig. LABEL:fig_x1 depicts the state evolution of subsystem 1 (similar results for other subsystems are omitted). It can be seen that the variance of the system state trajectories under Algorithm 2 is much larger than those under Algorithm 4. In addition, Fig. LABEL:fig_u shows the evolution of the global constraint. It can be found that there exist constraint violations in Algorithm 2. However, owing to the implementation scheme developed in Section IV, our approach can guarantee the satisfaction of the global constraint.
(a) Evolution of (a) subsystem 1 and (b) global constraint.